# MEEG and AT-DGNN: Improving EEG Emotion Recognition with Music Introducing and Graph-based Learning

### v3

@@ Abstract
We present the MEEG dataset, a multi-modal collection of music-induced electroencephalogram (EEG) recordings designed to capture emotional responses to various musical stimuli across different valence and arousal levels. This public dataset facilitates an in-depth examination of brainwave patterns within musical contexts, providing a robust foundation for studying brain network topology during emotional processing. Leveraging the MEEG dataset, we introduce the Attention-based Temporal Learner with Dynamic Graph Neural Network (AT-DGNN), a novel framework for EEG-based emotion recognition. This model combines an attention mechanism with a dynamic graph neural network (DGNN) to capture intricate EEG dynamics. The AT-DGNN achieves state-of-the-art (SOTA) performance with an accuracy of 83.74% in arousal recognition and 86.01% in valence recognition, outperforming existing SOTA methods. Comparative analysis with traditional datasets, such as DEAP, further validates the model’s effectiveness and underscores the potency of music as an emotional stimulus. This study advances graph-based learning methodology in brain-computer interfaces (BCI), significantly improving the accuracy of EEG-based emotion recognition. The MEEG dataset and source code are publicly available at https://github.com/xmh1011/AT-DGNN.
@@ I Introduction
Emotion, a complex physiological and psychological phenomenon, plays a critical role in interpreting genuine reactions to daily interactions [1]. In 1980, Russell introduced the dimensional models of emotion, with the Valence-Arousal (VA) model being particularly noteworthy [2]. The model conceptualize emotions within a continuous dimensional space, providing a framework that allows for a more nuanced quantification of emotional states.
BCI technology decodes EEG signals, which reflect cortical voltage fluctuations, into commands for external devices, enhancing human-computer interaction [3]. This technology involves preprocessing, feature extraction, and classification stages, translating thoughts into actionable commands. Due to their objectivity, high temporal resolution, affordability, and rapid acquisition, EEG signals are indispensable for emotion recognition and neuroscience research [4, 5].
In 2011, Koelstra et al. established the DEAP dataset, which consists of physiological signals, such as EEG, elicited by emotional responses to musical videos [6]. Subsequently, Zheng et al. introduced the SEED dataset in 2015, comprising well-annotated EEG signals induced by cinematic stimuli, with data collected from multiple subjects [7]. These datasets have significantly contributed to the advancement of computational models for emotion analysis. Over the past decades, a variety of machine learning and signal processing techniques have been applied to emotion recognition based on physiological data.
@@
All the aforementioned research treated EEG signals primarily as two-dimensional time-series data, with dimensions corresponding to channels and time [12]. Here, channels denote EEG electrodes, positioned on the scalp using the 10-20 system, as prescribed by the International Federation of Clinical Neurophysiology, to capture neural activity from distinct brain regions [3]. More recently, a paradigm shift has occurred with EEG data now being conceptualized as graphs, representing the electrodes’ two-dimensional arrangement on the scalp [13]. In this model, raw data or extracted features from each electrode augment this spatial framework by adding a third dimension [14].
In 2023, Ding et al. introduced the Local-Global-Graph Network (LGGNet) for processing EEG data in an image-like format, which improved classification performance [15]. Alternatively, LGGNet’s single-layer graph neural network faces challenges in capturing the dynamic features of EEG signals. Additionally, widely used EEG datasets such as DEAP and SEED primarily involve visual stimuli, which limits exploration of the relationship between music, emotions, and EEG signals. Conversely, EEG datasets and research using musical stimuli are relatively scarce. Moreover, existing SOTA methods achieve lower accuracy on the DEAP dataset, thus limiting comprehensive performance evaluation of models.
@@ II-A Multi-head Attention Mechanism
EEG data is inherently complex, characterized by temporal dependencies and discrete features. Utilizing multiple attention heads allows us to capture a diverse range of temporal patterns and dependencies, effectively identifying key temporal features. In this module, the attention block consists of a multi-head attention (MHA) layer with several self-attention heads.
@@ II-B Graph Neural Networks
II-C Graph Neural Networks for EEG
In 2019, Song et al. [17] first introduced a novel DGCNN specifically for multichannel EEG emotion recognition, employing an adjacency matrix to dynamically model EEG channel relationships and enhance feature discrimination. This method proved to be superior to existing approaches. Subsequently, advancements in GNNs have been increasingly applied to emotion recognition. On the SEED dataset, Bao et al. [18] integrated a multi-layer GNN with a style-reconfigurable CNN, while Asadzadeh et al. [19] further improved the DGCNN by incorporating Bayesian signal recovery techniques, both achieving enhanced performance. For the DEAP dataset, Zhao et al. [20] developed coherence-based node clustering and multi-pooling techniques for GCNs. Additionally, Zhang et al. [21] enhanced GCN feature extraction by using Granger causality analysis and differential entropy to construct more informative connectivity matrices.
@@ III-C2 Stacked DGNN
A stacked Directed Graph Neural Network (DGNN) captures complex graph relationships by recalculating the adjacency matrix at each layer based on input features, effectively capturing hierarchical feature variations.
$S=Z_{\textit{agg}}\cdot Z_{\textit{agg}}^{T}$ (16)
$A=S+I,\quad\tilde{A}=\tilde{D}^{-0.5}\cdot A\cdot\tilde{D}^{-0.5}$ (17)
Finally, the procedure of AT-DGNN can be summarized in Algorithm 1.
@@ IV-C Implement Details
TABLE I: Comparison of ACC and F1 scores (mean $\pm$ std) for the MEEG dataset under trial-wise $10$-fold cross-validation
Method Arousal Valence Emotion
ACC (%) std (%) F1 (%) std (%) ACC (%) std (%) F1 (%) std (%) ACC (%) F1 (%)
EEGNet 75.47 17.63 75.54 17.83 73.66 16.82 73.33 16.94 74.57 74.44
DeepConvNet 79.35 15.03 78.43 18.10 72.38 18.87 73.03 19.27 75.87 75.73
ShallowConvNet 80.96 14.49 80.87 14.71 79.77 16.44 80.22 15.60 80.37 80.55
TSception 79.23 11.51 81.30 9.84 78.34 15.03 81.18 12.08 78.79 81.24
EEG-TCNet 76.36 17.05 73.79 19.66 69.72 16.06 54.28 22.43 73.04 64.04
TCNet-Fusion 76.72 17.32 76.95 17.49 75.88 16.70 75.42 16.97 76.30 76,20
ATCNet 82.01 14.06 79.17 17.57 83.19 12.09 82.08 13.41 82.60 80.63
DGCNN 82.12 12.05 81.83 12.45 82.72 11.00 82.40 11.44 82.42 82.12
LGGNet-Fro 81.85 12.57 81.58 12.73 84.28 11.83 83.87 12.66 83.07 82.73
LGGNet-Gen 82.15 12.72 82.15 12.57 84.53 11.44 84.38 11.56 83.34 83.27
LGGNet-Hem 81.92 12.08 81.58 12.43 84.93 11.21 84.46 12.27 83.42 83.02
AT-DGNN-Fro 83.51 11.09 83.06 11.83 85.56 10.69 85.61 10.89 84.54 84.34
AT-DGNN-Gen 83.74 12.03 84.56 10.88 86.01 9.40 85.68 10.06 84.88 85.12
AT-DGNN-Hem 83.73 11.20 84.78 10.93 84.89 10.57 85.35 10.24 84.31 85.07
@@
  • The results for all models are derived from the optimal values among ten trials with distinct random seeds.
@@ V-A Emotion Recognition
As shown in Table I, the AT-DGNN series, particularly the AT-DGNN-Gen model, significantly outperforms the previous SOTA LGGNet series in both arousal and valence dimensions, with a notable increase in F1 scores and accuracy. The AT-DGNN-Gen model achieves the highest accuracy of 83.74% in the arousal dimension, surpassing the LGGNet-Gen model by 1.59%, and also records the highest accuracy of 86.01% in the valence dimension, marking a substantial improvement of 1.08% over the LGGNet series. This improvement underscores the efficacy of our approach in capturing complex patterns within the MEEG dataset, which is known for its challenging and class-imbalanced nature.
The AT-DGNN-Gen model exemplifies the superior performance of the AT-DGNN series, achieving robust results with an average accuracy of 84.88% and an F1 score of 85.12% across emotion dimensions. This underscores the effectiveness of our graph-based approach in handling complex emotion recognition tasks. Furthermore, consistent with the findings of LGGNet [15], the graph configuration $G_{g}$ demonstrated superior performance in both AT-DGNN and LGGNet models compared to other graph definitions, indicating its optimal structure for capturing relevant features in emotion analysis.
Furthermore, the results show a significant improvement over traditional CNN and TCN-based methods, with average gains in F1 scores and accuracy exceeding previous models by up to 7.16% and 5.78%, respectively. This highlights the enhanced capability of the DGNN approach in analyzing graph-structured data, offering more accurate and reliable emotion recognition. Additionally, the AT-DGNN exhibits a lower standard deviation compared to other models, indicating more balanced performance across various samples and superior generalization ability.
All methods achieve high accuracy on the MEEG dataset, with CNN and TCN models reaching 75% or higher, and GNN models exceeding 80%. On the DEAP dataset, AT-DGNN achieves an average accuracy of 60.56% across four dimensions, while LGGNet reaches 60.96%. According to Huang et al. [29], other SOTA methods also achieve around 60% accuracy on DEAP, with some slightly outperforming AT-DGNN and others not. All methods perform significantly better on the MEEG dataset, suggesting that music is more effective than video in evoking emotions.
@@ V-B Ablation Study
  • M: Multi-head attention mechanism.
  • D: Dynamic graph neural network.
@@
  • 1, 2, 3, 4: The number of stacked DGNN layers.
@@
As illustrated in Table II, an ablation study is conducted to systematically assess the individual contributions of key components within the AT-DGNN model: sliding window segmentation (S), multi-head attention mechanism (M), and dynamic graph neural network (D). The results underscore the pivotal roles of these components, indicating that their omission results in declines in accuracy (ACC) of 2.97%, 6.57%, and 3.99% and in F1 scores of 3.31%, 5.66%, and 3.96%, respectively, thereby demonstrating their functionality and synergy.
Furthermore, the experiment demonstrated that the number of layers in the DGNN significantly affects the model’s performance. Modifying the number of DGNN layers to $1$, $2$, or $4$ lead to reductions in accuracy of 2.97%, 1.94%, and 1.71%, and in F1 scores of 3.31%, 2.60%, and 1.64%, respectively. A configuration with a single layer implies non-utilization of the stacked DGNN, highlighting the importance of network depth tuning in optimizing the model’s capacity to process and analyze complex relationships in EEG data.
The optimal configuration, utilizing $3$ layers of the DGNN, demonstrated superior effectiveness by achieving the highest scores in both ACC and F1 metrics, substantiating the design choices and highlighting the critical interdependence of its three main components: sliding window segmentation, multi-head attention mechanisms, and dynamic graph neural networks. These findings reveal how the synergistic operation of these elements is essential for capturing the dynamic temporal features of EEG data, significantly enhancing the model’s capability to process EEG data effectively and advancing computational neuroscience analysis.
@@ VII ACKNOWLEDGEMENTS
This paper is supported in part by the Shenzhen Fundamental Research Program under Grant JCYJ20230807094104009.

### v4

@@ Abstract
We present the MEEG dataset, a multi-modal collection of music-induced electroencephalogram (EEG) recordings designed to capture emotional responses to various musical stimuli across different valence and arousal levels. This public dataset facilitates an in-depth examination of brainwave patterns within musical contexts, providing a robust foundation for studying brain network topology during emotional processing. Leveraging the MEEG dataset, we introduce the Attention-based Temporal Learner with Dynamic Graph Neural Network (AT-DGNN), a novel framework for EEG-based emotion recognition. This model combines an attention mechanism with a dynamic graph neural network (DGNN) to capture intricate EEG dynamics. The AT-DGNN achieves state-of-the-art (SOTA) performance with an accuracy of 83.74% in arousal recognition and 86.01% in valence recognition, outperforming existing SOTA methods. This study advances graph-based learning methodology in brain-computer interfaces (BCI), significantly improving the accuracy of EEG-based emotion recognition. The MEEG dataset and source code are publicly available at https://github.com/xmh1011/AT-DGNN.
@@ I Introduction
Emotion plays a key role in interpreting human responses, and Russell’s Valence-Arousal (VA) model offers a dimensional framework for quantifying emotional states within a continuous space [1]. Meanwhile, BCI technology, which decodes EEG signals reflecting cortical voltage fluctuations, translates thoughts into commands for external devices, significantly enhancing human-computer interaction [2]. With their high temporal resolution, objectivity, affordability, and quick acquisition, EEG signals are crucial for emotion recognition and neuroscience research, particularly in integrating emotional models with advanced decoding methods [3].
In 2011, Koelstra et al. introduced the DEAP dataset, consisting of physiological signals like EEG elicited by emotional responses to musical videos [4], followed by Zheng et al.’s SEED dataset in 2015, which provided well-annotated EEG signals from cinematic stimuli across multiple subjects [5]. These datasets have greatly advanced computational models for emotion analysis.
@@
All aforementioned studies treated EEG signals as two-dimensional time-series data, with channels representing EEG electrodes positioned according to the 10-20 system to capture neural activity across brain regions [10, 2]. Recently, EEG data have been increasingly modeled as graphs, representing the spatial arrangement of electrodes, where features from each electrode add a third dimension to the model [11, 12].
In 2019, Song et al. introduced a novel DGCNN for multichannel EEG emotion recognition, utilizing an adjacency matrix to dynamically model EEG channel relationships and enhance feature discrimination, demonstrating superiority over existing approaches [13]. Subsequently, advancements in GNNs have been increasingly applied to emotion recognition. Bao et al. integrated a multi-layer GNN with a style-reconfigurable CNN [14], while Asadzadeh et al. further improved the DGCNN by incorporating Bayesian signal recovery techniques, both achieving enhanced performance [15].
Building on prior advancements, greater progress was made in 2023 when Ding et al. introduced the Local-Global-Graph Network (LGGNet) for processing EEG data in an image-like format, achieving improved classification performance [16]. Alternatively, LGGNet’s single-layer graph neural network faces challenges in capturing the dynamic features of EEG signals. Additionally, widely used EEG datasets such as DEAP and SEED primarily involve visual stimuli, which limits exploration of the relationship between music, emotions, and EEG signals. Conversely, EEG datasets and research using musical stimuli are relatively scarce. Moreover, existing SOTA methods achieve lower accuracy on the DEAP dataset, thus limiting comprehensive performance evaluation of models.
@@ II-A Multi-head Attention Mechanism
EEG data is inherently complex, with significant temporal dependencies and discrete features. Multi-head attention (MHA) helps capture diverse temporal patterns and dependencies, enabling the identification of key features. In this module, the attention block consists of a multi-head attention layer with several self-attention heads. As shown by Zhang et al. [17], multi-head self-attention effectively integrates multi-source information, improving the model’s robustness in handling complex temporal features in EEG data, leading to enhanced classification performance.
@@ III-C2 Stacked DGNN
A stacked Dynamic Graph Neural Network (DGNN) captures complex graph relationships by recalculating the adjacency matrix at each layer based on input features, effectively capturing hierarchical feature variations.
$S=Z_{\textit{agg}}\cdot Z_{\textit{agg}}^{T}$ (16)
$A=S+I,\quad\tilde{A}=\tilde{D}^{-0.5}\cdot A\cdot\tilde{D}^{-0.5}$ (17)
Finally, the procedure of AT-DGNN can be summarized in Algorithm 1.
@@ IV-C Implement Details
TABLE I: Comparison of ACC and F1 scores for the MEEG dataset and DEAP dataset under trial-wise 10-fold cross-validation
Method MEEG Arousal MEEG Valence MEEG Emotion DEAP Emotion
ACC (%) F1 (%) ACC (%) F1 (%) ACC (%) F1 (%) ACC (%) F1 (%)
EEGNet 75.47∗∗ 75.54∗∗ 73.66∗∗ 73.33∗∗ 74.57 74.44 60.36 64.64
DeepConvNet 79.35∗∗ 78.43∗∗ 72.38∗∗ 73.03∗∗ 75.87 75.73 60.41 63.80
ShallowConvNet 80.96∗∗ 80.87∗∗ 79.77∗∗ 80.22∗∗ 80.37 80.55 61.01 64.44
TSception 79.23∗∗ 81.30 78.34∗∗ 81.18∗∗ 78.79 81.24 60.42 64.15
EEG-TCNet 76.36∗∗ 73.79∗∗ 69.72∗∗ 54.28∗∗∗ 73.04 64.04 56.11 52.50
TCNet-Fusion 76.72∗∗ 76.95∗∗ 75.88∗∗ 75.42∗∗ 76.30 76.20 59.33 63.94
ATCNet 82.01∗∗ 79.17∗∗ 83.19∗∗ 82.08∗∗ 82.60 80.63 60.83 62.32
DGCNN 82.12∗∗ 81.83∗∗ 82.72∗∗ 82.40∗∗ 82.42 82.12 60.36 61.27
LGGNet-Fro 81.85∗∗ 81.58∗∗ 84.28∗∗ 83.87∗∗ 83.07 82.73 60.59 64.48
LGGNet-Gen 82.15∗∗ 82.15∗∗ 84.53∗∗ 84.38∗∗ 83.34 83.27 60.96 65.41
LGGNet-Hem 81.92∗∗ 81.58∗∗ 84.93∗∗ 84.46∗∗ 83.42 83.02 59.88 62.50
AT-DGNN-Fro 83.51∗∗ 83.06∗∗ 85.56∗∗ 85.61∗∗ 84.54 84.34 59.95 62.61
AT-DGNN-Gen 83.74∗∗ 84.56∗∗ 86.01 85.68 84.88 85.12 60.55 63.83
AT-DGNN-Hem 83.73∗∗ 84.78∗∗ 84.89∗∗ 85.35∗∗ 84.31 85.07 59.36 62.16
  • * indicating $std<10\%$, ** indicating $std<20\%$, *** indicating $std<30\%$.
@@
  • The results for are derived from the optimal values among ten trials with distinct random seeds.
@@ V-A Emotion Recognition
As shown in Table I, the AT-DGNN series, particularly the AT-DGNN-Gen model, significantly outperforms the previous state-of-the-art LGGNet series in both arousal and valence dimensions, with notable increases in F1 scores and accuracy. Specifically, the AT-DGNN-Gen model achieves the highest accuracy of 83.74% in the arousal dimension, surpassing the LGGNet-Gen model by 1.59%, and records the highest accuracy of 86.01% in the valence dimension, marking a substantial improvement of 1.08% over the LGGNet series. This improvement underscores the efficacy of our approach in capturing complex patterns within the MEEG dataset, which is known for its challenging and class-imbalanced nature.
Moreover, the AT-DGNN-Gen model exemplifies the superior performance of the AT-DGNN series, achieving robust results with an average accuracy of 84.88% and an F1 score of 85.12% across emotion dimensions. This highlights the effectiveness of our graph-based approach in handling complex emotion recognition tasks. Furthermore, consistent with the findings of LGGNet [16], the graph configuration $G_{g}$ demonstrated superior performance in both AT-DGNN and LGGNet models compared to other graph definitions, indicating its optimal structure for capturing relevant features in emotion analysis.
The results also show a significant improvement over traditional CNN and TCN-based methods, with average gains in F1 scores and accuracy exceeding previous models by up to 7.16% and 5.78%, respectively. This highlights the enhanced capability of the DGNN approach in analyzing graph-structured data, offering more accurate and reliable emotion recognition. Additionally, the AT-DGNN exhibits a lower standard deviation compared to other models, indicating more balanced performance across various samples and superior generalization ability.
On both the MEEG and DEAP datasets, the AT-DGNN demonstrates strong generalization ability. All models achieve high accuracy on the MEEG dataset, with CNN and TCN models reaching 75% or higher, and GNN models exceeding 80%. Specifically, all models, including AT-DGNN, achieve significantly higher accuracy on the MEEG dataset than on the DEAP dataset, suggesting that music is more effective than video in evoking emotions and indicating the higher quality of the MEEG dataset. On the DEAP dataset, the AT-DGNN achieves an average accuracy of 60.56% across four dimensions, close to other SOTA models and surpassing some of them, which demonstrates the AT-DGNN model’s strong generalization capability across different datasets.
@@ V-B Ablation Study
  • A: Multi-head attention mechanism.
  • G: Graph neural network.
  • T: Temporal learner.
@@
  • 1, 2, 3, 4: The number of DGNN layers or temporal learner.
@@
As shown in Table II, the ablation study systematically evaluates the contributions of key components in the AT-DGNN-Gen model, including sliding window segmentation (S), multi-head attention mechanism (A), graph neural network layers (G), and temporal learner layers (T). The baseline model, incorporating all components with three layers each for G and T, achieved the highest ACC of 84.88% and F1 score of 85.12%.
Excluding the multi-head attention mechanism (A) caused a significant performance drop, with ACC decreasing by 6.57% and F1 score by 5.66%. Similarly, removing the sliding window segmentation (S) resulted in reductions of 3.99% in ACC and 3.96% in F1 score. These results confirm the essential roles of S and A in feature extraction and temporal modeling.
Varying the number of G layers showed that using three layers yielded optimal results, while deviations caused performance declines. Increasing G layers to four resulted in a minor ACC drop of 1.71%, while reducing to one layer led to a larger decline of 2.97%. Similar trends were observed for T layers, where three layers provided the best balance, and exclusion of T caused a substantial drop of 10.24% in ACC and 10.28% in F1 score, highlighting its critical role in capturing temporal dynamics.
Overall, the ablation study demonstrates that each component of the AT-DGNN-Gen model contributes significantly to its performance. The sliding window segmentation and multi-head attention mechanism are crucial for effective feature extraction, while the optimal number of layers for the graph neural network and temporal learner is three, balancing model complexity and performance. These findings substantiate the design choices of our model and underscore the importance of the synergistic operation of its components in enhancing emotion recognition from EEG data. The results emphasize how the model’s architecture is essential for capturing the complex spatiotemporal patterns inherent in EEG signals, thereby advancing computational neuroscience analysis.
@@ VII ACKNOWLEDGEMENTS
This paper is supported by the Shenzhen Fundamental Research Program under Grant JCYJ20230807094104009. and Key Lab of Information Network Security, Ministry of Public Security.